OpenStudy (anonymous):

Emily went to a toy store that was selling one Nintendo DS game and one Wii game for a total of $70. Emily bought 3 DS games and 2 Wii games and spent $160. What was the price of one Wii game

OpenStudy (anonymous):

let Nintendo's price be n and Wii's price be w as per question 1n+1w=$70......equation 1 and 3n+2w=$160.....equation 2 now multiply equation 1 by 2 making it 2n+2w=$140 now subtract it from equation 2 and we get n=$20 put it back in equation 1 and we get w=$(70-20)=$50